# Clairaut anti-invariant Riemannian map from/to Kähler manifolds admitting Ricci solitons

### v1

@@ 4 Clairaut Anti-invariant Riemannian maps to Kähler manifolds admitting Ricci Solitons
$rangeF_{*}=span\{F_{*}{X_{1}}=e_{1}^{\prime},F_{*}{X_{2}}=e_{3}^{\prime}\}$ and $(rangeF_{*})^{\perp}=\{e_{2}^{\prime},e_{4}^{\prime},e_{5}^{\prime},e_{6}^{\prime}\}.$
Also we can see that $JF_{*}{X_{1}}=e_{2}^{\prime},JF_{*}{X_{2}}=e_{4}^{\prime}$ means $F$ is an anti-invariant Riemannian map.
@@
$\Gamma^{1}_{14}=-1=\Gamma^{1}_{41},\Gamma^{3}_{34}=-1=\Gamma^{3}_{43},\Gamma^{4}_{11}=e^{-2y_{4}},\Gamma^{4}_{33}=e^{-2y_{4}}$ and rest of Christoffel symbols are zero.
@@
$\nabla_{F_{*}{X_{1}}}{F_{*}{X_{1}}}=e_{4}^{\prime},\nabla_{F_{*}{X_{1}}}e_{4}^{\prime}=-e_{1}^{\prime},\nabla_{F_{*}{X_{2}}}{F_{*}{X_{2}}}=e_{4}^{\prime},\nabla_{F_{*}{X_{2}}}e_{4}^{\prime}=-e_{3}^{\prime}$ and rest of covariant derivatives are zero.
@@
$\nabla^{M}g=\sum_{i,j=1}^{6}g^{ij}\frac{\partial g}{\partial y_{i}}\frac{\partial}{\partial y_{j}}.$
Therefore, we get $\nabla^{M}g=-e_{4}^{\prime}.$
Since $TM=rangeF_{*}\oplus(rangeF_{*})^{\perp},$ for any $T_{i}=a_{i}F_{*}X_{1}+b_{i}e_{2}^{\prime}+c_{i}F_{*}X_{2}+d_{i}e_{4}^{\prime}+h_{i}e_{5}^{\prime}+l_{i}e_{6}^{\prime},i=1,...,6.$
Since
$(L_{T_{1}}g_{M})(T_{2},T_{3})=g_{M}(\nabla_{T_{2}}T_{1},T_{3})+g_{M}(\nabla_{T_{3}}T_{1},T_{2}).$
Substituting the values of covariant derivatives, we obtain
$\begin{split}\frac{1}{2}(L_{T_{1}}g_{M})(T_{2},T_{3})=\frac{1}{2}\Big{(}a_{1}a_{2}d_{3}-a_{2}d_{1}a_{3}+c_{1}c_{2}d_{3}-c_{2}c_{3}d_{1}+a_{1}a_{3}d_{2}-a_{3}d_{1}a_{2}+c_{3}c_{1}d_{2}-c_{3}d_{1}c_{2}\Big{)},\end{split}$
$g_{M}(T_{2},T_{3})=(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+h_{2}h_{3}+l_{2}l_{3}),$
$Ric(T_{2},T_{3})=-a_{2}a_{3}-d_{2}d_{3}.$
Let $(M,g_{M},J)$ admits a Ricci soliton, then
$\frac{1}{2}(L_{Y_{1}}g_{M})(Y_{2},Y_{3})+Ric(Y_{2},Y_{3})+\lambda g_{M}(Y_{2},Y_{3})=0$
which implies
$\lambda=\frac{a_{2}a_{3}+d_{2}d_{3}+a_{2}a_{3}d_{1}+c_{2}c_{3}d_{1}-\frac{1}{2}\Big{(}a_{1}a_{2}d_{3}+c_{1}c_{2}d_{3}+a_{1}a_{3}d_{2}+c_{3}c_{1}d_{2}\Big{)}}{(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+h_{2}h_{3}+l_{2}l_{3})},$
where $a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+h_{2}h_{3}+l_{2}l_{3}\neq 0.$ For the suitable choice of scalar’s viz, $a_{i},b_{i},c_{,}d_{i},h_{i},l_{i}$ the behaviour of Ricci soliton will be expanding, steady, shrinking according to $\lambda>0,$ $\lambda=0,$ $\lambda<0$.
Example 4.2.
Let $M=\{(x_{1},x_{2},x_{3},x_{4}):x_{i}\neq 0,\forall i=1,...,4\}$ be a Riemannian manifold with Riemannian metric $g_{M}=dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2}+dx_{4}^{2}$ and $N=R^{4}$ be a Kähler manifold with Riemannian metric $g_{N}=e^{-2y_{3}}dy_{1}^{2}+e^{-2y_{3}}dy_{2}^{2}+dy_{3}^{2}+e^{-2y_{3}}dy_{4}$ and complex structure $J(a,b,c,d)=(-d,-c,-b,-a).$ Now we construct a map $F:(M,g_{M})\rightarrow(N,g_{N},J),$ defined by
$F(x_{1},x_{2},x_{3},x_{4})=\Big{(}\frac{x_{1}+x_{2}}{\sqrt{2}},\frac{x_{3}+x_{4}}{\sqrt{2}},0,0\Big{)}.$
After some calculations, we get
$kerF_{*}=span\Big{\{}U_{1}=\frac{e_{1}-e_{2}}{\sqrt{2}},U_{2}=\frac{e_{3}-e_{4}}{\sqrt{2}}\Big{\}}$
and
$(kerF_{*})^{\perp}=span\Big{\{}X_{1}=\frac{e_{1}+e_{2}}{\sqrt{2}},X_{2}=\frac{e_{3}+e_{4}}{\sqrt{2}}\Big{\}},$
where $\{e_{1}=\frac{\partial}{\partial x_{1}},e_{2}=\frac{\partial}{\partial x_{2}},e_{3}=\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}}\}$ and $\{e_{1}^{\prime}=e^{y_{3}}\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=e^{y_{3}}\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=e^{y_{3}}\frac{\partial}{\partial y_{4}}\}$ are the bases of $T_{p}M$ and $T_{F(p)N}$ respectively.
Here, $F_{*}(X_{1})=e_{1}^{\prime},F_{*}(X_{2})=e_{2}^{\prime}$ and we can easily compute that
$g_{M}(X_{i},X_{i})=g_{N}(F_{*}{X_{i}},F_{*}{X_{i}})~{}~{}\forall X_{i}\in\Gamma(TM).$
$rangeF_{*}=span\{F_{*}(X_{1})=e_{1}^{\prime},F_{*}(X_{2})=e_{2}^{\prime}\}$ and $(rangeF_{*})^{\perp}=\{e_{3}^{\prime},e_{4}^{\prime}\}.$
Also we can see that $JF_{*}(X_{1})=-e_{4}^{\prime},JF_{*}{X_{2}}=-e_{3}^{\prime},$ means $F$ is a Langrarian Riemannian map.
To show that $F$ is a Clairaut Langragian Riemanian map, we have to find $g$ such that
$(\nabla F_{*})(X,X)=-g(X,X)\nabla^{M}g.$
After some computation, we get that for Riemannian metric $g_{M},$ Christoffel symbols are given by
$\Gamma^{1}_{13}=-1=\Gamma^{1}_{31},\Gamma^{2}_{23}=-1=\Gamma^{2}_{32},\Gamma^{3}_{11}=e^{-2y_{3}},\Gamma^{3}_{22}=e^{-2y_{3}},\Gamma^{3}_{44}=e^{-2y_{3}},\Gamma^{4}_{34}=-1=\Gamma^{4}_{43}$ and rest of Christoffel symbols are zero.
The covariant derivative are calculated as
$\nabla_{F_{*}{X_{1}}}F_{*}{X_{1}}=e_{3}^{\prime},\nabla_{F_{*}{X_{1}}}e_{3}^{\prime}=-e_{1}^{\prime},\nabla_{F_{*}{X_{2}}}F_{*}{X_{2}}=e_{3}^{\prime},\nabla_{F_{*}{X_{2}}}e_{3}^{\prime}=-e_{2}^{\prime},\nabla_{e_{4}^{\prime}}e_{3}^{\prime}=-e_{4}^{\prime},\nabla_{e_{4}^{\prime}}e_{4}^{\prime}=e_{3}^{\prime}$and rest of covariant derivatives are zero.
Gradient of any smooth function $g$ is given by
$\nabla^{M}g=\sum_{i,j=1}^{4}g^{ij}\frac{\partial g}{\partial y_{i}}\frac{\partial}{\partial y_{j}}.$
Therefore, we get $\nabla^{M}g=-e_{3}^{\prime}.$
After some calculations, using (3.1), (3.2), (3.3) and above covariant derivatives, we get
$Ric(F_{*}X_{1},F_{*}X_{1})=-2,Ric(e_{3}^{\prime},e_{3}^{\prime})=-2,Ric(e_{4}^{\prime},e_{4}^{\prime})=-2$ and rest are zero.
Since $TM=rangeF_{*}\oplus(rangeF_{*})^{\perp},$ for any $T_{i}=a_{i}F_{*}X_{1}+b_{i}F_{*}X_{2}+c_{i}e_{3}^{\prime}+d_{i}e_{4}^{\prime},i=1,...,4.$
Since
$(L_{T_{1}}g_{M})(T_{2},T_{3})=g_{M}(\nabla_{T_{2}}T_{1},T_{3})+g_{M}(\nabla_{T_{3}}T_{1},T_{2}).$
Substituting the values of covariant derivatives in the above equation, we obtain
$\begin{split}\frac{1}{2}(L_{T_{1}}g_{M})(T_{2},T_{3})=&\frac{1}{2}\Big{(}a_{1}a_{2}c_{3}+b_{1}b_{2}c_{3}+d_{1}d_{2}c_{3}+a_{1}a_{3}c_{2}+b_{1}b_{3}c_{2}+d_{3}d_{1}c_{2}\Big{)}\\ &-a_{2}c_{1}a_{3}-b_{2}b_{3}c_{1}-d_{2}d_{3}c_{1},\end{split}$
$Ric(T_{2},T_{3})=-2a_{2}a_{3}-2c_{2}c_{3}-2d_{2}d_{3},$
and
$g_{M}(T_{2},T_{3})=(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}).$
We know that $(M,g_{M},J)$ admits a Ricci soliton, if
$\frac{1}{2}(L_{Y_{1}}g_{M})(Y_{2},Y_{3})+Ric(Y_{2},Y_{3})+\lambda g_{M}(Y_{2},Y_{3})=0$
this implies
$\lambda=\frac{P_{1}-P_{2}}{a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}},a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}\neq 0.$
where
$P_{1}=a_{2}a_{3}(c_{1}+2)+d_{2}d_{3}(c_{1}+2)+b_{2}b_{3}c_{1}+2c_{2}c_{3},$
$P_{2}=\frac{1}{2}\Big{(}a_{1}a_{2}c_{3}+b_{1}b_{2}c_{3}+d_{1}d_{2}c_{3}+a_{1}a_{3}c_{2}+b_{1}b_{3}c_{2}+d_{3}d_{1}c_{2}\Big{)},$
for the suitable choice of scalar’s viz, $a_{i},b_{i},c_{,}d_{i}\in\operatorname{\mathbb{R}},$ the behaviour of Ricci soliton will be expanding, steady, shrinking according to $\lambda>0,$ $\lambda=0,$ $\lambda<0$.
5 Acknowledgement
@@
First author is thankful to the Department of Science and Technology(DST) Government of India for providing financial assistance in terms of FIST project(TPN-69301) vide the letter with Ref No.:(SR/FST/MS-1/2021/104). The second author is grateful to the financial support provided by CSIR (Council of science and industrial research) Delhi, India. File no.[09/1051(12062)/2021-EMR-I].

### v2

@@ 4 Clairaut Anti-invariant Riemannian maps to Kähler manifolds admitting Ricci Solitons
$rangeF_{*}=span\{F_{*}{X_{1}}=e_{2}^{\prime},F_{*}{X_{2}}=e_{5}^{\prime}\}$ and $(rangeF_{*})^{\perp}=\{e_{1}^{\prime},e_{3}^{\prime},e_{4}^{\prime},e_{6}^{\prime}\}.$
Also we can see that $J^{\prime}F_{*}{X_{1}}=-e_{1}^{\prime},J^{\prime}F_{*}{X_{2}}=e_{6}^{\prime}$ which implies that $F$ is an anti-invariant Riemannian map.
@@
$\Gamma^{3}_{35}=1=\Gamma^{3}_{53},\Gamma^{6}_{56}=1=\Gamma^{6}_{65},\Gamma^{5}_{33}=-e^{2y_{5}},\Gamma^{5}_{66}=-e^{2y_{5}}$ and rest of Christoffel symbols are zero.
@@
$\nabla_{e_{3}^{\prime}}{e_{5}^{\prime}}=e_{3}^{\prime},\nabla_{e_{6}^{\prime}}e_{5}^{\prime}=e_{6}^{\prime},\nabla_{e_{3}^{\prime}}{e_{3}^{\prime}}=-e_{5}^{\prime},\nabla_{e_{6}^{\prime}}e_{6}^{\prime}=-e_{5}^{\prime}$ and rest of covariant derivatives are zero.
Since, $(\nabla F_{*})(X,X)\in\Gamma(rangeF_{*})^{\perp}$ for any $X\in\Gamma(kerF_{*})^{\perp}.$ So, here we can write $(\nabla F_{*})(X,X)=\eta_{1}e_{1}^{\prime}+\eta_{2}e_{3}^{\prime}+\eta_{3}e_{4}^{\prime}+\eta_{4}e_{6}^{\prime},$ and $g_{M}(X,X)=g_{M}(\delta_{1}X_{1}+\delta_{2}X_{2},\delta_{1}X_{1}+\delta_{2}X_{2})=\delta_{1}^{2}+\delta_{2}^{2}=\delta^{2}$ for some $\eta_{1},\eta_{2},\delta_{1},\delta_{2}\in\mathbb{R}.$
@@
$\nabla^{N}g=\sum_{i,j=1}^{6}g^{ij}\frac{\partial g}{\partial y_{i}}\frac{\partial}{\partial y_{j}}.$
Therefore
$\nabla^{N}g=\frac{\partial g}{\partial y_{1}}\frac{\partial}{\partial y_{1}}+\frac{\partial g}{\partial y_{2}}\frac{\partial}{\partial y_{2}}+e^{-2y_{5}}\frac{\partial g}{\partial y_{3}}\frac{\partial}{\partial y_{3}}+\frac{\partial g}{\partial y_{4}}\frac{\partial}{\partial y_{4}}+\frac{\partial g}{\partial y_{5}}\frac{\partial}{\partial y_{5}}+e^{-2y_{5}}\frac{\partial g}{\partial y_{6}}\frac{\partial}{\partial y_{6}}.$
Thus
$\nabla^{N}g=\frac{\eta_{1}}{\delta^{2}}e_{1}^{\prime}+e^{-y_{5}}\frac{\eta_{2}}{\delta^{2}}e_{3}^{\prime}+\frac{\eta_{3}}{\delta^{2}}e_{4}^{\prime}+e^{-y_{5}}\frac{\eta_{4}}{\delta^{2}}e_{6}^{\prime}$
for the function $g=\frac{\eta_{1}}{\delta^{2}}y_{1}+e^{-y_{5}}\frac{\eta_{2}}{\delta^{2}}y_{3}+\frac{\eta_{3}}{\delta^{2}}y_{4}+e^{-y_{5}}\frac{\eta_{4}}{\delta^{2}}y_{6}.$ Further, proceeding as in example 3.1, one can easily show that $N$ admits a Ricci soliton.
5 Conclusion and future scope
We have successfully investigated the geometry of Clairaut anti-invariant Riemannian maps with Kähler and Ricci soliton structures. In the future, one can study Clairaut invariant [33], semi-invariant [9, 19] and slant [34] Riemannian maps with Kähler and Ricci soliton structures.
Acknowledgement
@@
First author is grateful to the financial support provided by CSIR (Council of science and industrial research) Delhi, India. File no.[09/1051(12062)/2021-EMR-I]. The second author is thankful to the Department of Science and Technology(DST) Government of India for providing financial assistance in terms of FIST project(TPN-69301) vide the letter with Ref No.:(SR/FST/MS-1/2021/104).